# Dynamical patterns and nonreciprocal effective interactions in an active-passive mixture through exact hydrodynamic analysis

### v2

@@ Abstract
11footnotetext: bruna@maths.ox.ac.uk22footnotetext: These authors contributed equally to this work.
@@ Results
Figure 1: Phase diagrams spanned by the active and passive particles’ volume fractions $\phi_{a}$ and $\phi_{p}$, respectively, for (a) $\text{Pe}=5$ and (b) $\text{Pe}=7.5$. The spinodal (blue curve) encloses the region of linear stability of homogeneous solutions. The binodal (red curve) encloses the region of phase separation. Intersection points of the tie lines (gray lines) with the binodal indicate the composition of the liquid and vapor phases. Black crosses indicate bifurcations of co-dimension two; see text for discussion. Regions with stable Homogeneous (H), Phase Separated (PS), and Dynamic (D) steady states are marked on the diagram. [In (b), the white region is left unlabelled because the boundary between H+D and H is unknown and the grey region corresponds to $\phi_{a}+\phi_{p}>1$, which is inaccessible due to size exclusion. The black tie line intersects the spinodal, so phase-separated states above this line are linearly unstable.]
We consider an active-passive lattice gas (APLG) model that extends the active lattice gases of [24, 19, 25]. It is defined on a two-dimensional periodic square lattice with spacing $h$. Placing at most one particle per site, we populate the lattice with three types of particles $\sigma\in\{+1,-1,0\}$: active particles oriented right ($\sigma=+1$), active particles oriented left ($\sigma=-1$) and passive particles ($\sigma=0$). Each lattice site $(i,j)$ has position $\bm{x}=(ih,jh)\in[0,\ell_{x})\times[0,\ell_{y})$. The model dynamics can be split into four parts: (i) passive particles attempt nearest neighbour jumps with jump rate $D_{T}/h^{2}$ per adjacent site, where $D_{T}$ is the spatial diffusion constant; (ii) active particles perform nearest neighbor random walk, weakly biased in the direction of their orientation to account for self-propulsion. In particular, a jump in the $\bm{u}$ direction (where $|\bm{u}|=h$) is attempted at rate $D_{T}/h^{2}+\frac{v_{0}}{2h^{2}}(\bm{u}\cdot{{{\bm{e}_{\sigma}}}})$, where ${{{\bm{e}_{\sigma}}}}=(\sigma,0)$ is the particle’s orientation, and $v_{0}$ is the self-propulsion speed; (iii) both types of particles are under an exclusion rule: if the target site of a jump is occupied, the jump is aborted; if the site is otherwise empty, the jump is executed; and (iv) each active particle orientation flips at rate $D_{R}$. The total numbers of active and passive particles are specified via their volume fractions $\phi_{a}$ and $\phi_{p}$, respectively, and the overall volume fraction is $\phi=\phi_{a}+\phi_{p}\in[0,1]$.
@@
To take the hydrodynamic limit, it is convenient to rescale time and space by $D_{R}^{-1}$ and $\sqrt{D_{T}/D_{R}}$ respectively, and introduce the Péclet number $\text{Pe}=v_{0}/\sqrt{D_{T}D_{R}}$. A configuration of the APLG is defined in terms of occupancies: $\eta_{\sigma}(\bm{x},t)\in\{0,1\}$ is the number of particles of type $\sigma$ at site $\bm{x}$ and time $t$. The hydrodynamic limit equations describe the evolution of the local densities $\rho_{\sigma}(\bm{x},t)$ of particles of type $\sigma\in\{+1,-1,0\}$, as the lattice spacing $h\to 0$. Building on [19], we rigorously derive the hydrodynamic limit of the APLG model, obtaining exact macroscopic evolution equations for the densities $\rho_{\sigma}$ (see Appendix A)
@@
[…]
@@
with $\alpha=\pi/2-1$, which exactly matches the value and the first derivative of $d_{s}(\rho)$ at $\rho=0,1$, and approximates $d_{s}$ very accurately for all $\rho$. For other accurate approximations, see [30, 31, 32, 33, 34, 35]. Here, we use (3) in the numerical analysis of the exact hydrodynamic limit (1). Corrections to the mean-field approximation $\langle\eta_{\sigma}(\bm{x},t)\eta_{\sigma}(\hat{\bm{x}},t)\rangle\approx\rho_{\sigma}(\bm{x},t)\rho_{\sigma}(\hat{\bm{x}},t)$ are important in the APLG model and enter the hydrodynamic limit; the mean-field approximation would lead to $d_{s}(\rho)=(1-\rho)$. The method to obtain (1) is valid for any dimension greater than one.
@@
Figure 2: (a),(b) Kymographs showing the spatiotemporal dynamics of a phase-separating solution. (a) $y$-averaged local density of a particle simulation (see Methods). (b) Numerical solution to (1), the initial condition is a homogenous state with a uniform random perturbation (see Appendix 1 for details, Eq. (S62)). (c) Density profile at $t=20$. Black lines display the solution to (1), and red lines display the local density of the particle simulation. Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.5$, $\phi_{p}=0.1$, $\Delta x=0.01$ (PDE model), $h=0.01$ (particle model).
Figure 3: (a),(b) Kymographs showing the spatiotemporal dynamics of a CP solution. (a) The $y$-averaged local density of a particle simulation (see Methods). (b) The solution to (1) whose initial condition is a homogenous state with a uniform random perturbation (see Appendix 1, Eq. (S62)). (c) Density profiles before, during, and after the collision of two counter-propagating interfaces ($t=17,18,19,20$). Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.3$, $\phi_{p}=0.6$, $\Delta x=0.01$, $h=0.01$.
@@
Figure 4: Traveling profile solution to (1), whose initial condition is the homogeneous state perturbed by a combination of left and right traveling unstable sinusoidal modes and a uniform random perturbation (see Appendix 1, Eqs. (S62,S63) for further details). (a) Density profile at $t=500$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=25$, $\phi_{a}=0.36$, $\phi_{p}=0.3$, $\Delta x=0.05$. Speed of propagation in (6) is $c=-1.8832$.
@@
Figure 5: T solutions of the finite-$L$ problem (7) and the large-$L$ problem (8) at $\text{Pe}=7.5$. (Top) Solutions to (8) as a function of $z/L$. (Bottom) Solutions plotted parametrically on the phase diagram of Fig. 1(b), with dashed lines for solutions of (7) and solid lines for solutions of (8). Volume fractions $(\phi_{a},\phi_{p})$ for each column are given in the top, also marked as black triangles in the bottom row. Other parameters: $N=1024$ for numerical solution of (8) and $L=25$, $N=1600$ for (7). Wave speeds for (a,b,c,d) are $c=1.95,1.39,0.41,0.31$ (2 decimal places).
@@
Recall that $L\gg 1$ corresponds to $\ell_{x}\gg\sqrt{D_{T}/D_{R}}$: this means that the system size is much larger than the intrinsic length scale associated with particle motion. Some active-matter and nonreciprocal models [9, 48] support traveling phase-separated states where a large system has macroscopic liquid and vapor domains separated by sharp interfaces, which move at constant velocity. However, this is not possible in the APLG. To see this, substitute (6) into (1) and write $z=x-ct/L$ to obtain
$-(c/L)\varrho_{\sigma}^{\prime}=\partial_{z}\left[d_{s}(\varrho)\varrho_{\sigma}^{\prime}+\varrho_{\sigma}{\cal D}(\varrho)\varrho^{\prime}\right]\\ -\text{Pe}\partial_{z}\left[\varrho_{\sigma}s(\varrho)\mathrm{m}+\sigma d_{s}(\varrho)\varrho_{\sigma}\right]-\sigma\mathrm{m},$ (7)
@@
There are macroscopically smooth T solutions of (7) satisfying $\varrho^{\prime}_{\sigma}=O(1/L)$ [with $c=O(1)$ and $\mathrm{m}=O(1/L)$]. We do find such solutions numerically (see Methods). However, Fig. 4(a) includes an interfacial region where $\varrho$ varies quickly in space, hinting at the existence of solutions with traveling narrow interfaces [$\varrho^{\prime}=O(1)$].
Figure 6: T solutions of the finite-$L$ problem (7) for different values of $L$ and the large-$L$ problem (8) at $\text{Pe}=7.5$. Total density $\varrho$ (top) and magnetisation $\mathrm{m}$ (bottom). The magnetisation for $L=\infty$ is $\mathrm{m}=0$ for $|z|>0$. Parameters as in Fig. 5(b).
Figure 7: Phase diagrams from Fig. 1(b), overlayed with the types of long-time solutions of (1): homogeneous (H), phase-separated (PS), traveling (T), and counter-propagating (CP), see Methods for the classification. The initial condition is the homogeneous state, perturbed by (a) left and right traveling unstable sinusoidal modes (b) left traveling unstable sinusoidal modes (see Appendix 1 for details). Parameters: $\text{Pe}=7.5$, $L=25.0$, $\phi_{a}=0.36$, $\phi_{p}=0.3$, $\Delta x=0.05$.
@@
$-(c/L)\varrho_{a}^{\prime}=$ $\partial_{z}\left[d_{s}(\varrho)\varrho_{a}^{\prime}+\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}\right]$ (8a)
@@
$-(c/L)\varrho_{0}^{\prime}=$ $\partial_{z}\left[d_{s}(\varrho)\varrho_{0}^{\prime}+\varrho_{0}{\cal D}(\varrho)\varrho^{\prime}\right]$ (8b)
@@
The array of solutions in Fig. 5 illustrates the rich phenomenology of the APLG in large domains. In particular, the appearance of narrow interfaces in T solutions is a surprising feature, since it connects these dynamically-patterned states to the equilibrium-like constructions of the binodal and the associated interfacial profiles. While the analytical characterization of T solutions is not possible for CP solutions, numerical simulations show that CP solutions also feature narrow interfaces that follow the tie lines (see Appendix, Fig. S5). The distinction between narrow interfaces and macroscopically smooth profiles would be very challenging to characterize by direct numerical solution of (1): the method of matched asymptotics makes this possible.
@@ Discussion
We introduced the APLG as a microscopic model of interacting particles and characterized its hydrodynamic behavior. In addition to PS states familiar from pure active systems, we find rich behavior, including that the spinodal curve can protrude through the binodal. This signals the existence of dynamical steady states, which we classify according to their symmetries. Some of these results are similar to previous work on the NRCH equation [10, 9, 11, 13], but our hydrodynamic PDE includes the magnetization $m$ as a slow hydrodynamic variable, in addition to the two conserved densities $\rho_{a},\rho_{0}$; it also has a distinct set of nonlinearities. Hence, our approach of deriving hydrodynamic equations exactly from a microscopic model complements the generic description of nonreciprocal systems via the NRCH equation. As noted in Appendix .3, it is important that the APLG supports the complex behavior of other nonreciprocal models, despite the idealized modeling assumptions in its definition.
@@ Methods
$-\epsilon^{2}c\bar{\varrho}_{\sigma}^{\prime}=\epsilon^{2}\partial_{\bar{z}}\left[d_{s}(\bar{\varrho})\bar{\varrho}_{\sigma}^{\prime}+\bar{\varrho}_{\sigma}{\cal D}(\bar{\varrho})\bar{\varrho}^{\prime}\right]\\ -\epsilon\text{Pe}\partial_{\bar{z}}\left[\bar{\varrho}_{\sigma}s(\bar{\varrho})\bar{\mathrm{m}}+\sigma d_{s}(\bar{\varrho})\bar{\varrho}_{\sigma}\right]-\sigma\bar{\mathrm{m}},$ (14)
@@
$-c\bar{\varrho}_{\sigma}^{(0)\prime}=\partial_{\bar{z}}\left[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(0)\prime}+\bar{\varrho}_{\sigma}^{(0)}{\cal D}(\bar{\varrho}^{(0)})\bar{\varrho}^{(0)\prime}\right]-\sigma\bar{\mathrm{m}}^{(2)}\\ -\text{Pe}\partial_{\bar{z}}\!\left[\bar{\varrho}_{\sigma}^{(0)}s(\bar{\varrho}^{(0)})\bar{\mathrm{m}}^{(1)}\!+\!\sigma d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(1)}\!+\!\sigma d_{s}^{\prime}(\bar{\varrho}^{(0)})\bar{\varrho}^{(1)}\bar{\varrho}_{\sigma}^{(0)}\right],$ (15)
@@
$-c\bar{\varrho}_{a}^{(0)\prime}=\partial_{\bar{z}}\left[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{a}^{(0)\prime}+\bar{\varrho}_{a}^{(0)}{\cal D}(\bar{\varrho}^{(0)})\bar{\varrho}^{(0)\prime}\right]\\ -\text{Pe}\partial_{\bar{z}}\left[\left(\bar{\varrho}_{a}^{(0)}s(\bar{\varrho}^{(0)})+d_{s}(\bar{\varrho}^{(0)})\right)\bar{\mathrm{m}}^{(1)}\right],$ (16)
@@
    $\sup_{t\in\mathcal{T}^{*}}d_{\text{H}}(t)<0.05$to H solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$ and $\sup_{t\in\mathcal{T}^{*}}\tilde{c}(t)<0.01$to PS solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in\mathcal{T}^{*}}|\tilde{c}^{\prime}(t)|<10^{-5}$to T solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in\mathcal{T}^{*}}|\tilde{c}^{\prime}(t)|\geq 10^{-5}$to CP solution.
$\bullet$ Traveling profiles: The profiles $\varrho_{\sigma}(z)$ and speed $c$ in Fig. 5 are obtained numerically by discretizing Eqs. (7,8) with second-order centered differences and solving the resulting zeroth-finding problem subject to mass constraints $\int\varrho_{0}\mathrm{d}z=\phi_{p}$ and $\int\varrho_{+}\mathrm{d}z=\int\varrho_{-}\mathrm{d}z=\phi_{a}/2$ in $[-L/2,L/2]$. The finite-$L$ system (7) is initialized with a preexisting T solution with similar parameters or a steady state of (1) and solved subject to periodic boundary conditions and $\varrho(0)=\phi$ without loss of generality (since the problem has translational symmetry). For $L\gg 1$, the numerical procedure to solve (8) is as follows. Given parameter values $\phi$ and $\nu$ (5), we determine the inner solution via the abovementioned Coexisting Phases procedure. This results in liquid $(\varrho_{\sigma})_{l}$ and vapor $(\varrho_{\sigma})_{v}$ values, which become boundary values for the outer problem (8). [If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means that there is no inner region and we may proceed to solve (8) subject to periodic boundary conditions as in the finite $L$ case.] Then (8) is discretized using second-order finite-differences and solved in $[0,L]$ subject to Dirichlet boundary conditions $\varrho_{\sigma}(0)=(\varrho_{\sigma})_{v}$ and $\varrho_{\sigma}(L)=(\varrho_{\sigma})_{l}$ and mass constraints as above. If no solution is found, it indicates there may be a second interface. We initialize with a previous single interface T solution with similar parameters. We then place a new interface at $z=\text{argmax}_{z}|\partial_{z}\varrho|$. In both finite $L$ and $L\gg 1$ cases, we solve the resulting systems of equations using the NonlinearSolve.jl package in Julia (see Appendix 2).
@@
The particle and PDE simulations data that support the findings of this study have been deposited in the figshare repository [58].
@@
The code that supports the findings within this paper is available at the GitHub repository [59].

### v3

@@ Results
Figure 1: Phase diagrams of the active-passive mixture for two values of Pe. We show the diagram spanned by the active and passive particles’ volume fractions $\phi_{a}$ and $\phi_{p}$, respectively, for (a) $\text{Pe}=5$ and (b) $\text{Pe}=7.5$. The spinodal (blue curve) encloses the region of linear stability of homogeneous solutions. The binodal (red curve) encloses the region of phase separation. Intersection points of the tie lines (gray lines) with the binodal indicate the composition of the liquid and vapor phases. Black crosses indicate bifurcations of co-dimension two; see text for discussion. Regions with stable Homogeneous (H), Phase Separated (PS), and Dynamic (D) steady states are marked on the diagram. [In (b), the white region is left unlabelled because the boundary between H+D and H is unknown and the grey region corresponds to $\phi_{a}+\phi_{p}>1$, which is inaccessible due to size exclusion. The black tie line intersects the spinodal, so phase-separated states above this line are linearly unstable.]
We consider an active-passive lattice gas (APLG) model that extends the active lattice gases of [24, 19, 25]. It is defined on a two-dimensional periodic square lattice with spacing $h$. Placing at most one particle per site, we populate the lattice with three types of particles $\sigma\in\{+1,-1,0\}$: active particles oriented right ($\sigma=+1$), active particles oriented left ($\sigma=-1$) and passive particles ($\sigma=0$). Each lattice site $(i,j)$ has position ${\mathbf{x}}=(ih,jh)\in[0,\ell_{x})\times[0,\ell_{y})$. The model dynamics can be split into four parts: (i) passive particles attempt nearest neighbour jumps with jump rate $D_{T}/h^{2}$ per adjacent site, where $D_{T}$ is the spatial diffusion constant; (ii) active particles perform nearest neighbor random walk, weakly biased in the direction of their orientation to account for self-propulsion. In particular, a jump in the $\mathbf{u}$} direction (where $|{\mathbf{u}}|=h$) is attempted at rate $D_{T}/h^{2}+\frac{v_{0}}{2h^{2}}({\mathbf{u}}\cdot{\mathbf{e}}_{\sigma})$, where ${\mathbf{e}}_{\sigma}=(\sigma,0)$ is the particle’s orientation, and $v_{0}$ is the self-propulsion speed; (iii) both types of particles are under an exclusion rule: if the target site of a jump is occupied, the jump is aborted; if the site is otherwise empty, the jump is executed; and (iv) each active particle orientation flips at rate $D_{R}$. The total numbers of active and passive particles are specified via their volume fractions $\phi_{a}$ and $\phi_{p}$, respectively, and the overall volume fraction is $\phi=\phi_{a}+\phi_{p}\in[0,1]$.
@@
To take the hydrodynamic limit, it is convenient to rescale time and space by $D_{R}^{-1}$ and $\sqrt{D_{T}/D_{R}}$ respectively, and introduce the Péclet number $\text{Pe}=v_{0}/\sqrt{D_{T}D_{R}}$. A configuration of the APLG is defined in terms of occupancies: $\eta_{\sigma}({\mathbf{x}},t)\in\{0,1\}$ is the number of particles of type $\sigma$ at site $\mathbf{x}$} and time $t$. The hydrodynamic limit equations describe the evolution of the local densities $\rho_{\sigma}({\mathbf{x}},t)$ of particles of type $\sigma\in\{+1,-1,0\}$, as the lattice spacing $h\to 0$. Building on [19], we rigorously derive the hydrodynamic limit of the APLG model, obtaining exact macroscopic evolution equations for the densities $\rho_{\sigma}$ (see Appendix A)
@@
Without activity ($\text{Pe}=0)$, the APLG model reduces to a three-species symmetric simple exclusion process (SSEP), so it is natural that $d_{s}$ appears in its hydrodynamic limit [26]. Technically, the APLG model is of nongradient type in the sense of [18]; this means that while (1) is exact, there is no explicit expression for $d_{s}(\rho)$, although variational formulae are available [27, 28]. Even so, $d_{s}$ may be approximated to arbitrary accuracy, either numerically or analytically. We use the polynomial approximation [29]:
@@
with $\alpha=\pi/2-1$, which exactly matches the value and the first derivative of $d_{s}(\rho)$ at $\rho=0,1$, and approximates $d_{s}$ very accurately for all $\rho$. For other accurate approximations, see [30, 31, 32, 33, 34, 35]. Here, we use (3) in the numerical analysis of the exact hydrodynamic limit (1). Corrections to the mean-field approximation $\langle\eta_{\sigma}({\mathbf{x}},t)\eta_{\sigma}(\hat{\mathbf{x}},t)\rangle\approx\rho_{\sigma}({\mathbf{x}},t)\rho_{\sigma}(\hat{\mathbf{x}},t)$ are important in the APLG model and enter the hydrodynamic limit; the mean-field approximation would lead to $d_{s}(\rho)=(1-\rho)$. The method to obtain (1) is valid for any dimension greater than one.
@@
Figure 2: Phase-separated stationary solution. (a),(b) KymographsAppendix showing the spatiotemporal dynamics of a phase-separating solution. (a) $y$-averaged local density of a particle simulation (see Methods). (b) Numerical solution to (1), the initial condition is a homogenous state with a uniform random perturbation (see Appendix 1 for details, Eq. (S62)). (c) Density profile at $t=20$. Black lines display the solution to (1), and red lines display the local density of the particle simulation. Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.5$, $\phi_{p}=0.1$, $\Delta x=0.01$ (PDE model), $h=0.01$ (particle model).
Figure 3: Counter-propagating solution (a),(b) Kymographs showing the spatiotemporal dynamics of a CP solution. (a) The $y$-averaged local density of a particle simulation (see Methods). (b) The solution to (1) whose initial condition is a homogenous state with a uniform random perturbation (see Appendix 1, Eq. (S62)). (c) Density profiles before, during, and after the collision of two counter-propagating interfaces ($t=17,18,19,20$). Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.3$, $\phi_{p}=0.6$, $\Delta x=0.01$, $h=0.01$.
@@
Figure 4: Traveling solution. Traveling profile solution to (1), whose initial condition is the homogeneous state perturbed by a combination of left and right traveling unstable sinusoidal modes and a uniform random perturbation (see Appendix 1, Eqs. (S62,S63) for further details). (a) Density profile at $t=500$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=25$, $\phi_{a}=0.36$, $\phi_{p}=0.3$, $\Delta x=0.05$. Speed of propagation in (6) is c=1.8832c=-1.8832.
@@
Figure 5: Traveling solutions in the large system-size limit. T solutions of the finite-$L$ problem (7) and the large-$L$ problem (8) at $\text{Pe}=7.5$. (Top) Solutions to (8) as a function of $z/L$. (Bottom) Solutions plotted parametrically on the phase diagram of Fig. 1(b), with dashed lines for solutions of (7) and solid lines for solutions of (8). Volume fractions $(\phi_{a},\phi_{p})$ for each column are given in the top, also marked as black triangles in the bottom row. Other parameters: $N=1024$ for numerical solution of (8) and $L=25$, $N=1600$ for (7). Wave speeds for (a,b,c,d) are $c=1.95,1.39,0.41,0.31$ (2 decimal places).
@@
Recall that $L\gg 1$ corresponds to $\ell_{x}\gg\sqrt{D_{T}/D_{R}}$: this means that the system size is much larger than the intrinsic length scale associated with particle motion. Some active-matter and nonreciprocal models [9, 48] support traveling phase-separated states where a large system has macroscopic liquid and vapor domains separated by sharp interfaces, which move at constant velocity. However, this is not possible in the APLG. To see this, substitute (6) into (1) and write z=xct/Lz=x-ct/L to obtain
$(c/L)\varrho_{\sigma}^{\prime}=\partial_{z}\left[d_{s}(\varrho)\varrho_{\sigma}^{\prime}+\varrho_{\sigma}{\cal D}(\varrho)\varrho^{\prime}\right]\\ -\text{Pe}\partial_{z}\left[\varrho_{\sigma}s(\varrho)\mathrm{m}+\sigma d_{s}(\varrho)\varrho_{\sigma}\right]-\sigma\mathrm{m},$ (7)
@@
There are macroscopically smooth T solutions of (7) satisfying $\varrho^{\prime}_{\sigma}=O(1/L)$ [with $c=O(1)$ and $\mathrm{m}=O(1/L)$]. We do find such solutions numerically (see Methods). However, Fig. 4(a) includes an interfacial region where $\varrho$ varies quickly in space, hinting at the existence of solutions with traveling narrow interfaces $\varrho^{\prime}=O(1)$].
Figure 6: Domain-size dependence of traveling solution. T solutions of the finite-$L$ problem (7) for different values of $L$ and the large-$L$ problem (8) at $\text{Pe}=7.5$. Total density $\varrho$ (top) and magnetisation $\mathrm{m}$ (bottom). The magnetisation for $L=\infty$ is $\mathrm{m}=0$ for $|z|>0$. Parameters as in Fig. 5(b).
Figure 7: Stability of T and CP solutions. Phase diagrams from Fig. 1(b), overlayed with the types of long-time solutions of (1): homogeneous (H), phase-separated (PS), traveling (T), and counter-propagating (CP), see Methods for the classification. The initial condition is the homogeneous state, perturbed by (a) left and right traveling unstable sinusoidal modes (b) left traveling unstable sinusoidal modes (see Appendix 1 for details). Parameters: $\text{Pe}=7.5$, $L=25.0$, $\phi_{a}=0.36$, $\phi_{p}=0.3$, $\Delta x=0.05$.
@@
$-(c/L)\varrho_{a}^{\prime}=\$$\partial_{z}\left[d_{s}(\varrho)\varrho_{a}^{\prime}+\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}\right]$ (8a)
@@
$-(c/L)\varrho_{0}^{\prime}=\$$\partial_{z}\left[d_{s}(\varrho)\varrho_{0}^{\prime}+\varrho_{0}{\cal D}(\varrho)\varrho^{\prime}\right]$ (8b)
@@
The array of solutions in Fig. 5 illustrates the rich phenomenology of the APLG in large domains. In particular, the appearance of narrow interfaces in T solutions is a surprising feature, since it connects these dynamically-patterned states to the equilibrium-like constructions of the binodal and the associated interfacial profiles. While the analytical characterization of T solutions is not possible for CP solutions, numerical simulations show that CP solutions also feature narrow interfaces that follow the tie lines (see Fig. S5, in Appendix). The distinction between narrow interfaces and macroscopically smooth profiles would be very challenging to characterize by direct numerical solution of (1): the method of matched asymptotics makes this possible.
@@ Discussion
We introduced the APLG as a microscopic model of interacting particles and characterized its hydrodynamic behavior. In addition to PS states familiar from pure active systems, we find rich behavior, including that the spinodal curve can protrude through the binodal. This signals the existence of dynamical steady states, which we classify according to their symmetries. Some of these results are similar to previous work on the NRCH equation [10, 9, 11, 13], but our hydrodynamic PDE includes the magnetization $m$ as a slow hydrodynamic variable, in addition to the two conserved densities $\rho_{a},\rho_{0}$; it also has a distinct set of nonlinearities. Hence, our approach of deriving hydrodynamic equations exactly from a microscopic model complements the generic description of nonreciprocal systems via the NRCH equation. As already noted in Sec. .3, it is important that the APLG supports the complex behavior of other nonreciprocal models, despite the idealized modeling assumptions in its definition.
@@ Methods
$\epsilon^{2}c\bar{\varrho}_{\sigma}^{\prime}=\epsilon^{2}\partial_{\bar{z}}\left[d_{s}(\bar{\varrho})\bar{\varrho}_{\sigma}^{\prime}+\bar{\varrho}_{\sigma}{\cal D}(\bar{\varrho})\bar{\varrho}^{\prime}\right]\\ -\epsilon\text{Pe}\partial_{\bar{z}}\left[\bar{\varrho}_{\sigma}s(\bar{\varrho})\bar{\mathrm{m}}+\sigma d_{s}(\bar{\varrho})\bar{\varrho}_{\sigma}\right]-\sigma\bar{\mathrm{m}},$ (14)
@@
$\hskip-10.0pt-c\bar{\varrho}_{\sigma}^{(0)\prime}=\partial_{\bar{z}}\left[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(0)\prime}+\bar{\varrho}_{\sigma}^{(0)}{\cal D}(\bar{\varrho}^{(0)})\bar{\varrho}^{(0)\prime}\right]-\sigma\bar{\mathrm{m}}^{(2)}\\ -\text{Pe}\partial_{\bar{z}}\!\left[\bar{\varrho}_{\sigma}^{(0)}s(\bar{\varrho}^{(0)})\bar{\mathrm{m}}^{(1)}\!+\!\sigma d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(1)}\!+\!\sigma d_{s}^{\prime}(\bar{\varrho}^{(0)})\bar{\varrho}^{(1)}\bar{\varrho}_{\sigma}^{(0)}\right],$ (15)
@@
$\hskip-10.0pt-c\bar{\varrho}_{a}^{(0)\prime}=\partial_{\bar{z}}\left[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{a}^{(0)\prime}+\bar{\varrho}_{a}^{(0)}{\cal D}(\bar{\varrho}^{(0)})\bar{\varrho}^{(0)\prime}\right]\\ -\text{Pe}\partial_{\bar{z}}\left[\left(\bar{\varrho}_{a}^{(0)}s(\bar{\varrho}^{(0)})+d_{s}(\bar{\varrho}^{(0)})\right)\bar{\mathrm{m}}^{(1)}\right],$ (16)
@@
    $\sup_{t\in\mathcal{T}^{*}}d_{\text{H}}(t)<0.05$ $\to$ H solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$ and $\sup_{t\in\mathcal{T}^{*}}\tilde{c}(t)<0.01$ $\to$ PS solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in\mathcal{T}^{*}}|\tilde{c}^{\prime}(t)|<10^{-5}$ $\to$ T solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in\mathcal{T}^{*}}|\tilde{c}^{\prime}(t)|\geq 10^{-5}$ $\to$ CP solution.
$\bullet$ Traveling profiles: The profiles $\varrho_{\sigma}(z)$ and speed $c$ in Fig. 5 are obtained numerically by discretizing Eqs. (7,8) with second-order centered differences and solving the resulting zeroth-finding problem subject to mass constraints $\int\varrho_{0}\mathrm{d}z=\phi_{p}$ and $\int\varrho_{+}\mathrm{d}z=\int\varrho_{-}\mathrm{d}z=\phi_{a}/2$ in [L/2,L/2][-L/2,L/2]. The finite-$L$ system (7) is initialized with a preexisting T solution with similar parameters or a steady state of (1) and solved subject to periodic boundary conditions and $\varrho(0)=\phi$ without loss of generality (since the problem has translational symmetry). For $L\gg 1$, the numerical procedure to solve (8) is as follows. Given parameter values $\phi$ and $\nu$ (5), we determine the inner solution via the abovementioned Coexisting Phases procedure. This results in liquid $(\varrho_{\sigma})_{l}$ and vapor $(\varrho_{\sigma})_{v}$ values, which become boundary values for the outer problem (8). [If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means that there is no inner region and we may proceed to solve (8) subject to periodic boundary conditions as in the finite $L$ case.] Then (8) is discretized using second-order finite-differences and solved in $[0,L]$ subject to Dirichlet boundary conditions $\varrho_{\sigma}(0)=(\varrho_{\sigma})_{v}$ and $\varrho_{\sigma}(L)=(\varrho_{\sigma})_{l}$ and mass constraints as above. If no solution is found, it indicates there may be a second interface. We initialize with a previous single interface T solution with similar parameters. We then place a new interface at $z=\text{argmax}_{z}|\partial_{z}\varrho|$. In both finite $L$ and $L\gg 1$ cases, we solve the resulting systems of equations using the NonlinearSolve.jl package in Julia (see Appendix 2).
@@
The simulation data generated in this study have been deposited in the Figshare database under accession code 10.25446/oxford.28881329 [58].
@@
[…]
Author contributions
JM, RLJ, and MB designed research, performed research, and wrote the paper.